# Stability of Non-Hyperbolic Sets

### v1

@@ 4.1 The combinatorics
Let $z^{(2)}=(z_{x}^{(2)},z_{y}^{(2)})=F_{t,a}^{n}(z^{(1)})$ and take a local unstable manifold $W^{u}_{\text{loc}}(z^{(2)})$ of vertical size $L\left(\lambda^{\theta}\mu\right)^{n}$ with $L$ a large constant which is going to be chosen later. Observe that $z^{(2)}$ is located close to $c$, see Figure 4.
The following lemma gives information on the size and the position of $W^{u}_{\text{\rm loc}}(z^{(2)})$. In particular Lemma 7 states how much $W^{u}_{\text{loc}}(z^{(2)})$ is shifted up with respect to the critical point $c$, see Figure 4. Its proof follows from Lemma 6 and from the fact that we are in the domain of linearization.
Lemma 7.
For $n$ large enough and $(x,y)\in W^{u}_{\text{\rm loc}}$, if $(t,a)\in\mathcal{B}_{n}$,
$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
$y-\mu_{0}^{n}z_{n,y}+\mu^{n}(a-a_{n}(t))$ $\leq$
$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right),$
where $z_{n}=z(t,a_{n}(t))$, $\mu_{0}=\mu(t,a_{n}(t))$ and $\mu=\mu(t,a)$. Moreover
$x=O\left(\lambda^{n}\right).$
In particular, when $a=a_{n}(t)$,
$y-\mu_{0}^{n}z_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$
The curves $W^{u}_{\text{\rm loc}}(z^{(1)})$, $W^{u}_{\text{\rm loc}}(z^{(2)})$ and $W^{u}_{\text{\rm loc}}(z^{(3)})$ (which will be introduced later) are graphs over smooth curves in the $x$-plane. We can describe these curves in coordinates centered at $z^{(i)}$. In particular, for each component of $W^{u}_{\text{\rm loc}}(z^{(i)})\setminus z^{(i)}$, there exists a smooth function $w$ such that if $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(i)})$ then $\Delta y=w\left(|\Delta x|\right)$.
In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(2)})$.
Lemma 8.
For $n\geq 1$ large enough, in a coordinate system centered in $z^{(2)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(2)})$ is given by
$\Delta y=Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|^{2}+O\left(|\Delta x|^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right),$
where $(\Delta x,\Delta y)\in W^{u}_{\text{\rm loc}}(z^{(2)})$, $0\leq\Delta y<L\left(\lambda^{\theta}\mu\right)^{n}$, $|\Delta x|=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$ and the constant $Q_{2}>0$ is a uniformly bounded and bounded away from zero.
@@
Consider the $W^{u}_{\text{loc}}(z^{(1)})$ of vertical size $L\lambda^{\theta n}$ and let $B\in T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})$ be a unit vector. Then in coordinate centered in $z^{(1)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{loc}}(z^{(1)})$ is given by
$\Delta y$ $=$ $Q_{1}|\Delta x|^{2}+O(|\Delta x|^{3}).$
where $Q_{1}$ is a uniform constant, see (4.5). Given the fact that $B$ has a non zero first coordinate, see $(F5)$, then the linear map $F^{n}_{t,a}$ turns the curve $W^{u}_{\text{loc}}(z^{(1)})$ into a curve of the form as stated in the lemma. Namely, recall that the curves $W^{u}_{\text{loc}}(z^{(1)})$ and $W^{u}_{\text{loc}}(z^{(2)})$ are graphs of functions over smooth curves in the $x$-plane. The curve $W^{u}_{\text{loc}}(z^{(1)})$ is parametrized by $|\Delta x|$ and by iteration by the linear map we obtain that the curve $W^{u}_{\text{loc}}(z^{(2)})$ is parametrized by $\lambda^{n}|\Delta x|$.
Observe also that for $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$, the highest vertical point of $W^{u}_{\text{loc}}(z^{(2)})$ is still in the domain of linearization. The factor $\left({\mu}/{\lambda^{2}}\right)^{n}$ comes from the change of curvature by $n$ linear iterates. ∎
Let $z^{(3)}=(z_{x}^{(3)},z_{y}^{(3)})$ be the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$. In the next lemma we prove that the new critical value $z^{(3)}$ is far from the critical value $z$ and from now on they will have independent behavior. In particular, while $z$ is contained in the basin of the sink, the new critical value $z^{(3)}$ it is outside of the basin and it will be involved in the creation of the new tangency.
Lemma 9.
@@
$0\leq Q(C-E)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)$
$\leq z_{y}^{(3)}(t,a)-z_{y}(t,a)\leq$
$Q(C+QL)^{2}\left(\lambda^{\theta}\mu\right)^{2n}+\frac{QT^{2}}{\mu^{n_{0}}}+O\left(\frac{\left(\lambda^{\theta}\mu\right)^{n}}{\mu^{n_{0}/2}}\right)+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
Moreover $T_{z^{(3)}}W^{u}_{\text{loc}}(z^{(3)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal. In particular,
$z_{y}^{(3)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))=QC^{2}\left(\lambda^{\theta}\mu\right)^{2n}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right).$
@@
Use coordinates centered at $c$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})=F_{t,a}^{n}\left(W^{u}_{\text{loc}}(z^{(1)})\right)$. Observe that
$\Delta y$ $=$ $y-\mu_{0}^{n}z_{n,y}+\mu_{0}^{n}z_{n,y}-1$ (4.6)
$=$ $y-\mu_{0}^{n}z_{n,y}+O\left(\lambda(t,a_{n}(t))|^{n}\right),$
where we used that $F^{n}_{t,a_{n}(t)}(z_{n})\in\Gamma$ which implies that
$|\mu_{0}^{n}z_{n,y}-1|=O\left(\left|F^{n}_{t,a_{n}(t)}(z_{n})_{x}\right|\right)=O\left(\lambda(t,a_{n}(t))^{n}\right).$
From Lemma 7 and for all $(t,a)\in\mathcal{B}_{n}$ we get
$(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right)$ $\leq\Delta y\leq$ (4.7)
$\frac{T}{\mu^{n_{0}/2}}+(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\lambda(t,a_{n}(t))^{n}\right),$ (4.8)
$|\Delta x|=O\left(\lambda^{n}\right).$ (4.9)
[…]
$\Delta\tilde{y}=O\left(\lambda^{n}\right)+Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right)=Q\left(\Delta y\right)^{2}+O\left(\left(\lambda^{\theta}\mu\right)^{3n}\right),$
[…]
In the following lemma we study the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$. One can observe that the curvature of $W^{u}_{\text{\rm loc}}(z^{(3)})$ has grown considerably compared to that of $W^{u}_{\text{\rm loc}}(z)$.
Lemma 10.
Let $(t,a)\in\mathcal{B}_{n}$ then $W^{u}_{\text{loc}}(z^{(3)}(t,a))$ is the graph of a function and its curvature satisfies
$\text{\rm curv}\left(W^{u}_{\text{loc}}(z^{(3)}(t,a))\right)\geq T\left(\frac{\mu^{4}}{\lambda^{2-3\theta}}\right)^{n},$
with $T>0$ a uniform constant and ${\mu^{4}}/{\lambda^{2-3\theta}}>1$.

### v3

@@ 4.1 The combinatorics
Let $(t,a)\in\mathcal{B}_{n}$ and let $z^{(2)}=(z_{x}^{(2)},z_{y}^{(2)})=F_{t,a}^{n}(z^{(1)})$ and take a local unstable manifold $W^{u}_{\text{loc}}(z^{(2)})$ of vertical size $L\left(\lambda^{\theta}\mu\right)^{n}$ with $L$ a large constant which is going to be chosen later, see Figure 4. The following lemma gives information on the size and the position of $W^{u}_{\text{\rm loc}}(z^{(2)})$.
Lemma 4.1.21.
For all $(t,a)\in\mathcal{B}_{n}$,
$z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right),$
where $z_{n}=z(t,a_{n}(t))$ and $\mu_{0}=\mu(t,a_{n}(t))$. Moreover if $(x,y)\in W^{u}_{\text{loc}}(z^{(2)})$,
$C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$
$\leq y-\mu_{0}^{n}z_{n,y}\leq$
$(C+QL)\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right),$
and
$x=\lambda^{n}+o\left(\lambda^{n}\right).$
@@
From Lemma 4.1.10, (2.2.9), (2.2.10), Remark 4.0.3 and from the fact that we are in the domain of linearization,
$z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}$ $=$ $z_{y}^{(2)}(t,a)-\mu^{n}z_{y}+\mu^{n}z_{y}-\mu^{n}_{0}z_{n,y}$
$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)+z_{n,y}\left(\mu^{n}-\mu^{n}_{0}\right)+\mu^{n}(a-a_{n}(t))$
$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)+O\left(\left(\frac{\mu}{\mu_{0}}\right)^{n}-1\right)+\mu^{n}(a-a_{n}(t))$
$=$ $C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right).$
Repeat the same estimates using the second inequalities from Lemma 4.1.10 to get the second inequalities in this lemma. For the last one, it is enough to use again Lemma 4.1.10 and the fact that we are in the linearization domain. ∎
The curve $W^{u}_{\text{\rm loc}}(z^{(2)})$, is a graph over the $y$-axes. In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(z^{(2)})$.
Lemma 4.1.22.
In a coordinate system centered in $z^{(2)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(2)})$ is given by
$\Delta y=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x^{2}+O\left(\Delta x^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right),$
where $0\leq\Delta y<L\left(\lambda^{\theta}\mu\right)^{n}$ and $\Delta x=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$.
Proof.
The lemma follows from the fact that the curve $W^{u}_{\text{loc}}(z^{(1)})$ is a graph over the $x$-axes and $W^{u}_{\text{loc}}(z^{(2)})$ is obtained by applying the linear map $F^{n}_{t,a}$ to the inequality defining the shape of $W^{u}_{\text{loc}}(z^{(1)})$ in Lemma 4.1.15. ∎
Let $z^{(3)}=(z_{x}^{(3)},z_{y}^{(3)})$ be the lowest point of $F^{N}_{t,a}\left(W^{u}_{\text{loc}}(z^{(2)})\right)$. At this moment we have three critical values, $z$, $z^{(1)}$ and $z^{(3)}$. The critical value $z^{(1)}$ is too closely connected to $z$ to use it to create a secondary tangency. In the next lemma we prove that the new critical value $z^{(3)}$ is far enough and independent enough from the critical value $z$ that we can use it to create the secondary tangency. This independent behavior of $z^{(3)}$ is the key of our constructions.
Lemma 4.1.23.
@@
$z_{y}^{(3)}(t,a)-z_{y}(t,a)=QH^{2}\left[1+O\left(H\right)\right].$
where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)$. In particular,
$z_{y}^{(3)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))=QC^{2}\left(\lambda^{\theta}\mu\right)^{2n}\left[1+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right].$
@@
Use coordinates $(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $c$. We are interested in the special point $m_{5}=(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(z^{(2)})$ such that $F^{N}(m_{5})=z^{(3)}$. In particular, if $F^{N}(m_{5})=(\Delta x_{z},\Delta y_{z})\in W^{u}_{\text{loc}}(z^{(3)})$ centered at $z$. Along the proof we will need as well coordinates $(\Delta x_{z^{(2)}},\Delta y_{z^{(2)}})\in W^{u}_{\text{loc}}(z^{(2)})$ centered at $z^{(2)}$. By construction and by Lemma 4.1.22 we get
$\Delta x_{c}=z^{(2)}_{x}-m_{5,x}+\Delta x_{z^{(2)}}=O\left(\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}\right)+\Delta x_{z^{(2)}},$ (4.1.24)
and
$\Delta y_{c}=z_{y}^{(2)}(t,a)-\mu^{n}_{0}z_{n,y}+\Delta y_{z^{(2)}}=H+\Delta y_{z^{(2)}}.$ (4.1.25)
Observe that, by Lemma 4.1.21 and by the definition of the strip $\mathcal{B}_{n}$, $H>\left(C-E\right)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left({\lambda^{\theta}}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)>O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$. Use in order, Lemma 2.2.12, (4.1.24), (4.1.25), (2.3.1), (2.3.2) and the fact that $\Delta x=O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$ (see Lemma 4.1.22) to get
$\Delta y_{z}$ $=$ $C\Delta x_{c}+Q\Delta y_{c}^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right)$
$=$ $C\Delta x_{z^{(2)}}+Q\Delta y_{z^{(2)}}^{2}+2QH\Delta y_{z^{(2)}}+QH^{2}+O\left(\lambda^{2+\theta}\right)^{\frac{n}{2}}$
$+$ $O\left(\Delta x_{z^{(2)}}^{2}+\Delta x_{z^{(2)}}H+\Delta x_{z^{(2)}}\Delta y_{z^{(2)}}+\Delta y_{z^{(2)}}^{3}+\Delta y_{z^{(2)}}^{2}H+\Delta y_{z^{(2)}}H^{2}+H^{3}\right)$
$=$ $C\Delta x_{z^{(2)}}\left[1+O(H)\right]+Q\Delta y_{z^{(2)}}^{2}\left[1+O(H)\right]+2QH\Delta y_{z^{(2)}}\left[1+O(H)\right]$
$+$ $QH^{2}\left[1+O(H)\right].$
Observe that the special point $m_{5}$ we are looking for will satisfy the equation $d\Delta y_{z}/d\Delta x_{z^{(2)}}=0$. Hence, by differentiating (4.1),
$0=\frac{d\Delta y_{z}}{d\Delta x_{z^{(2)}}}$ $=$ $C+2Q\Delta y_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}+2QH\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}$
$+$ $O\left(\Delta x_{z^{(2)}}+\Delta y_{z^{(2)}}+\Delta x_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{\Delta x_{z^{(2)}}}+\Delta y_{z^{(2)}}^{2}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}\right$.
$+$ $\left.\Delta y_{z^{(2)}}\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}H+\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}H^{2}+H\right).$
By Lemma 4.1.22,
$\Delta y_{z^{(2)}}$ $=$ $\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}+O\left(\Delta x_{z^{(2)}}^{3}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)$ (4.1.29)
$=$ $\frac{Q}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}^{2}\left[1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right]$
and
$\frac{d\Delta y_{z^{(2)}}}{d\Delta x_{z^{(2)}}}$ $=$ $2\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}+O\left(\Delta x_{z^{(2)}}^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)$
$=$ $2\frac{Q}{B^{2}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right].$
Using the last two inequalities in (4.1) we get
$0=\frac{d\Delta y_{z}}{d\Delta x_{z^{(2)}}}$ $=$ $2\frac{Q^{3}}{B^{4}}\left(\frac{\mu}{\lambda^{2}}\right)^{2n}\Delta x_{z^{(2)}}^{3}\left[1+O\left(H\right)\right]+4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(H\right)\right]$ (4.1.30)
$+$ $C\left[1+O(H)\right]$
where we also use Lemma 4.1.22 and (2.3.1). By (4.1.30) it follows that
$\Delta x_{z^{(2)}}=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right)$ (4.1.31)
and by (4.1.29),
$\Delta y_{z^{(2)}}=O\left(\left(\frac{\lambda^{2-2\theta}}{\mu^{3}}\right)^{n}\right).$ (4.1.32)
Use the last two inequalities in (4.1) to get
$\Delta y_{z}=QH^{2}\left[1+O(H)\right],$
where we also used the lower bound for $H$, (2.3.1) and (2.3.2). The lemma is proved. Moreover by using (4.1.31) in (4.1.30) we get an exact estimate for $\Delta x_{z^{(2)}}$. Namely,
$0$ $=$ $4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(2)}}\left[1+O\left(\left(\frac{\mu}{\lambda^{2}}\right)^{n}\frac{\Delta x_{z^{(2)}}^{2}}{H}\right)+O(H)\right]+C\left[1+O(H)\right]$
$=$ $4\frac{Q^{2}}{B^{2}}H\left(\frac{\mu}{\lambda^{2}}\right)^{2n}\Delta x_{z^{(2)}}\left[1+O(H)\right]+C\left[1+O(H)\right]$
and
$\Delta x_{z^{(2)}}=-\frac{CB^{2}}{4Q^{2}H}\left(\frac{\lambda^{2}}{\mu}\right)^{n}\left[1+O(H)\right].$ (4.1.33)
The curve $W^{u}_{\text{\rm loc}}(z^{(3)})$, is a graph over the $y$-axes. In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(z^{(3)})$.
Lemma 4.1.34.
In a coordinate system centered in $z^{(3)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(3)})$ is given by
$\Delta y=\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x^{2}+O\left(\Delta x^{3}\right),$
where $H=C\left(\lambda^{\theta}\mu\right)^{n}+\left(\mu^{n}+O(n)\right)\left(a-a_{n}(t)\right)$.